OpenStudy (anonymous):

HELP THIS IS A MATTER OF LIFE AND DEATH !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! !!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! !!!!!!!!!!!!!!! In triangle ABC, AB = x, BC = y, and CA = 2x. A similarity transformation with a scale factor of 0.5 maps triangle ABC to triangle MNO, such that vertices M, N, and O correspond to A, B, and C, respectively. If OM = 5, what is AB? Just trying to catch your attention since its my only way.

OpenStudy (anonymous):

A.AB = 2.5 B.AB = 10 C.AB = 5 D.AB = 1.25 E.AB = 2

OpenStudy (anonymous):

@TheSmartOne @eliassaab @AnswerMyQuestions @paki @satellite73 @dan815 @iGreen @SolomonZelman @linn99123

OpenStudy (anonymous):

Please help me out

OpenStudy (anonymous):

0_o

OpenStudy (anonymous):

I just need help understanding it. Give me steps to lead onto my answer please

OpenStudy (anonymous):

Please i just want to understand how to get to my answer

OpenStudy (anonymous):

all cant even help

OpenStudy (chosenmatt):

no i can help you :)

OpenStudy (anonymous):

@chosenmatt please thank you

OpenStudy (chosenmatt):

were applying a scale factor of 0.5 so

OpenStudy (anonymous):

i have drawn myself 2 triangles that represent ABC and MNO..

OpenStudy (chosenmatt):

mo=ac*0.5 mo=0.5*2x 5=0.5*2x so 5=x and x=5 :) so AB=5

OpenStudy (chosenmatt):

if you need anything else just tag me

TheSmartOne (thesmartone):



OpenStudy (chosenmatt):

mo=ac*0.5 mo=0.5*2x 5=0.5*2x so 5=x and x=5 :) so AB=5

TheSmartOne (thesmartone):



TheSmartOne (thesmartone):

Hope that helps you :D

OpenStudy (anonymous):

@chosenmatt Thank you :) but you did you get 5?@ thesmartone Thank you :)

OpenStudy (anonymous):

You gave the answer out. -.-

OpenStudy (anonymous):

nevermind i see now

TheSmartOne (thesmartone):

The answer was already given.. I made sure he did it correct...

OpenStudy (anonymous):

@AnswerMyQuestions atleast they explained it and decided to help me unlike you. I understand hw to do it now and ill use it for future problems like these

OpenStudy (chosenmatt):

your WELCOME ANYTIME can you medal someone please THANYOU

OpenStudy (anonymous):

>_>

OpenStudy (anonymous):

I wish i could metal you both for your time and help.

OpenStudy (anonymous):

medal*

jaynator495 (jaynator495):

I'm Answering Your Answered Questions ;3

jaynator495 (jaynator495):

life and death... well thats just rediculaus... LOL